OpenStudy (anonymous):

how do you make e^(ipi), e^(ipi)=-1

OpenStudy (kainui):

Well basically what you do is transform e^x, sin(x) and cos(x) into polynomials with infinite terms with a fancy thing called McLaurin series. Then you notice that e^x looks awfully similar to sine and cosine, and with a little manipulation you can see that: e^(ix)=cosx+isinx and from there you can plug in pi for x and find that e^ipi=-1 Here is an example of it, start there and watch the next video until you get to the final formula, it's fairly interesting! http://www.khanacademy.org/math/calculus/v/cosine-taylor-series-at-0--maclaurin

OpenStudy (anonymous):

I believe that this is a question involving complex nos. Remember that in complex numbers, e^(ix) = Cosx +iSin x So, e^(i pi) = Cos (pi) + i Sin (pi) = -1 +0 = -1

OpenStudy (anonymous):

ahh i see ty all